OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (jamesj):

Consider the equation x + 2 = 5 Saying that "x = 3 is a solution to the equation" means the value of x where x is 3 satisfies that equation, as indeed it does because 3 + 2 = 5.

OpenStudy (jamesj):

If you wanted to test whether x = 4 is a solution, you substitute it into the equation and see if the equation is satisfied. Well, if x = 4, then the equation x + 2 = 5 says that 4 + 2 = 5 But that is false. Hence x=4 is not a solution.